### 14 Mar 2026 v7

[…]
Gao Ming 20070602094@alu.cqu.edu.cn
(March 5, 2022)
[…]
In addition, this paper delves into the underlying mathematical laws of this new encoding mechanism, and develops a right multiplication operation to binary. Based on this right multiplication operation, we further constructed a nonlinear operation and designed another block symmetric encryption algorithm that is resistant to all forms of linear and differential attacks.
1 Introduction
Cryptography is one of the most important applications in the field of communication and computer science. In recent years, with the application of commerce, enterprises, banks and other departments, cryptography has been developed rapidly. Especially after Shannon put forward the mathematical analysis of security in “Communication theory of secrecy systems”[1], various design tools for cipher algorithms and corresponding attack tools have been developed one after another.
[…]
Among various attack methods, linear attack and differential attack are the two most core attack methods, which have become important analytical tools in the field of cryptography, driving the improvement of encryption algorithm.
Linear attack was first proposed by M. Matsui[2], this is an attack method that is currently applicable to almost all block encryption algorithms. Kaliski BS[3] proposed a multi-linear attack based on the linear attack, but the multi-linear attack has many limitations. And the Biryukov A[4] and Chao, J.Y[5] and others further improved the framework of multi-linear attacks, thus making linear attacks a larger application.
The differential attack method was first proposed by Eli Biham[6]. Biham E[7] extended it to a more powerful attack method. Tsunoo[8] further constructed multiple attack methods. These attack methods have extremely high skill in the attack process, which is worthy in-depth study.
In Chapter 3, we propose a novel encoding mechanism which we named Eagle. For the same key parameters, the Eagle encoding algorithm encodes the same input text sequence into completely different output text sequences by introducing randomness, and the output sequence can be considered completely random. The distinguishing feature of Eagle encoding algorithm from any other encoding algorithm is that for any given group of input and output texts, the key parameters are completely unsolvable, meaning that any key parameter can solve for the same input and output text.
[…]
In fact the encryption algorithm developed in Chapter 4 does not meet the security requirements for the key. This is because, when knowing the plaintext-ciphertext correspondence, although the complexity of cracking the key is exponential to the length of the key, it does not guarantee that only exhaustive search is effective for cracking the key. There may be some other attack methods that can effectively reduce the difficulty of cracking the key. Such as for a 256 bit key, although it can ensure that the computational complexity of cracking the key is at least $O(2^{64})$, this is still far from the ideal security of $O(2^{256})$. The encryption algorithm constructed in Chapter 4 only analyzes the computational complexity of cracking the key given a plaintext ciphertext correspondence, which does not guarantee its ability to resist linear and differential attacks, For attackers, they can construct plaintexts with certain characteristics multiple times, then find some structures that we did not anticipate from the relationship between plaintext and ciphertext.
In Chapter 5, we first delve into the mathematical laws behind the Eagle encoding algorithm and discover an operation that satisfies the allocation law for XOR operations. This operation is very similar to the allocation law of multiplication operations, but does not satisfy the commutative and associative laws, we refer to it as a “right multiplication operation on binary”. If we consider this right multiplication operation as a linear operation of one parameter on another, we can easily construct nonlinear operations of one parameter on another, such as “high-order” operations. Based on this nonlinear construction, we redesigned a block symmetric encryption algorithm and named it $Eagle^{*}$. For $Eagle^{*}$, knowing any plaintext ciphertext correspondence, there is no more effective way for attackers to crack the key than exhaustive search.
On the other hand, for the $Eagle^{*}$ encryption algorithm, the key is divided into three independent parts, and the ciphertext can also be seen as three independent parts. Two parts of the ciphertext is completely determined by the intermediate random number which is completely unrelated to the plaintext. The third part of the ciphertext is determined by the intermediate random number together with the plaintext sequence, but it is completely indistinguishable whether the uncertainty of the ciphertext brought by the random number or brought by different plaintexts. This theoretically ensure that $Eagle^{*}$ can resist all forms of linear attacks and differential attacks.
2 Notations
Some notations
$\oplus$ - XOR operation.
$D^{+n}$ - Do left cycle shift of D by n bits.
$m[i]$ - The $i$-th bit of $m$.
$\mathbb{E}_{w_{0},w_{1}}(s_{0},m)$ - Excute Eagle encoding algorithm as shown in Algorithm 1.
$\mathbb{D}_{w_{0},w_{1}}(c,s_{L})$ - Excute Eagle decoding algorithm as shown in Algorithm 2.
$m*\delta$ - Excute the right multiplication of $m$ with $\delta$ introduced in chapter 5.2.
$m*\delta^{2}$ - Excute the right multiplication of $m$ with $\delta$ twice as $m*\delta*\delta$.
$s/\delta$ - Excute the right division of $s$ to $\delta$ introduced in chapter 5.2.
$s/\delta^{2}$ - Excute the right division of $s$ to $\delta$ twice as $s/\delta/\delta$.
3 Introduction to Eagle encoding algorithm
We first introduce two common bit operations. XOR denoted as $\oplus$. Do left cycle shift of $D$ by $n$ bits which can be denoted as $D^{+n}$, for example $(10011010)^{+2}=(01101010)$ .
We select two $L$-bits parameters $w_{0}$ and $w_{1}$, have odd number of different bits. For example $w_{0}=10010011$ and $w_{1}=11000111$ have 3 bits (3 is odd) different.
$1\displaystyle w_{0}=1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}0{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}0{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}$1
$1\displaystyle w_{1}=1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}0{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}0{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}$1
We set the initial state of $L$-bit as $s_{0}$, we choose one parameter $w\in\{w_{0},w_{1}\}$, without loss of generality, assume that we choose $w=w_{0}$, then we define the following calculation
$s_{1}=w\oplus(s_{0}\oplus s_{0}^{+1})=w_{0}\oplus(s_{0}\oplus s_{0}^{+1})$ (3.1)
From (3.1), we can easily know
$s_{0}\oplus s_{0}^{+1}=s_{1}\oplus w$ (3.2)
If we only know $s_{1}$ but not $s_{0}$, we don ’t know whether $w=w_{0}$ or $w=w_{1}$ we used in (3.1), we can check it through a simple trial-and-error test. For example, we guess the parameter $w=w_{1}$ was used in (3.1), we need to find $s_{x}$ to satisfy
$s_{x}\oplus s_{x}^{+1}=s_{1}\oplus w_{1}$ (3.3)
In fact, since $w_{0}$ and $w_{1}$ have odd number of different bits, such $s_{x}$ does not exist. See Theorem 1 for details.
[Theorem 1] Two $L$-bit parameters $w_{0}$ and $w_{1}$ which have odd number of different bits, given $s_{0}$, we set $s_{1}=w_{0}\oplus(s_{0}\oplus s_{0}^{+1})$ , then there doesn’t exists $s_{x}$ satisfy $s_{x}\oplus s_{x}^{+1}=s_{1}\oplus w_{1}$.
Proof.
Firstly, by definition we have
$s_{1}\oplus w_{1}=w_{0}\oplus(s_{0}\oplus s_{0}^{+1})\oplus w_{1}=(s_{0}\oplus s_{0}^{+1})\oplus(w_{0}\oplus w_{1})$ (3.4)
Where $w_{0}$ and $w_{1}$ have odd number of different bits, so there are odd number of 1 in the bit string of $w_{0}\oplus w_{1}$ .
Proof by contradiction, we suppose that there exists $s_{x}$ satisfy $s_{x}\oplus s_{x}^{+1}=s_{1}\oplus w_{1}$, then by (3.4), we have
$s_{x}\oplus s_{x}^{+1}=(s_{0}\oplus s_{0}^{+1})\oplus(w_{0}\oplus w_{1})$ (3.5)
By simple calculation we have
$(s_{0}\oplus s_{x})\oplus(s_{0}\oplus s_{x})^{+1}=w_{0}\oplus w_{1}$ (3.6)
We set $s_{y}=s_{0}\oplus s_{x}$, then there are odd number of 1 in the bit string of $s_{y}\oplus s_{y}^{+1}$, without of generality, we suppose that the bits with 1 are $l_{1},l_{2},...,l_{u}$ ($u$ is odd). Compare to the first bit of $s_{y}$, the $l_{1}+1$ bit of $s_{y}$ is different from the first bit of $s_{y}$ , the $l_{2}+1$ bit of $s_{y}$ is the same with the first bit of $s_{y}$, the $l_{3}+1$ bit of $s_{y}$ is different from the first bit of $s_{y}$, the $l_{4}+1$ bit of $s_{y}$ is the same with the first bit of $s_{y}$, and so on, since $u$ is odd, u1u-1 is even, the $l_{u-1}+1$ bit of $s_{y}$ is the same with the first bit of $s_{y}$.
If $l_{u}<L$, the $l_{u}+1$ bit of $s_{y}$ is different from the first bit of $s_{y}$, so the last bit of $s_{y}$ is different from the first bit of $s_{y}$. On the other hand, the last bit of $s_{y}\oplus s_{y}^{+1}$ is 0, so the last bit of $s_{y}$ is the same with the first bit of $s_{y}$. This is contradictory.
[…]
If $l_{u}=L$, the last bit of $s_{y}$ is different from the the first bit of $s_{y}$. On the other hand, the last bit of $s_{y}$ is the same with $l_{u-1}$ bit of $s_{y}$ which is the same with the first bit of $s_{y}$. This is contradictory.
[…]
So there doesn’t exists $s_{x}$ satisfy $s_{x}\oplus s_{x}^{+1}=s_{1}\oplus w_{1}$ . ∎
Let’s go back to the discussion just now, after a trial-and-error test, we can accurately check which one ( $w=w_{0}$ or $w=w_{1}$ ) we just used in (3.1).
Now we suppose that there is a binary sequence $m=b_{1}b_{2}...b_{L}$ with length $L$. Start with $s_{0}$, read each bit of $M$ from left to right sequentially, when the bit $b[i](1\leq i\leq L)$ is 0, we set $s_{i}=w_{0}\oplus(s_{i-1}\oplus s_{i-1}^{+1})$, when the bit $b[i]$ is 1, we set $s_{i}=w_{1}\oplus(s_{i-1}\oplus s_{i-1}^{+1})$.
According to the above calculation, for every $s_{i}$, we can find the only $w_{x}=w_{0}$ or $w_{x}=w_{1}$ such that there exists $s_{i-1}$ satisfy $s_{i-1}\oplus s_{i-1}^{+1}=s_{i}\oplus w_{x}$. According to the properties of XOR and cyclic shift, we can easily see that there are only two $s_{i-1}$ that satisfy $s_{i-1}\oplus s_{i-1}^{+1}=s_{i}\oplus w_{x}$, and the two $s_{i-1}$ with each bit different. As long as we know any one bit of $s_{i-1}$, $s_{i-1}$ can be uniquely determined. So we only need to save one bit of $s_{i}$, finally by $s_{L}$, we can completely restore the original state $s_{0}$ and the binary sequence $m$.
Based on the above discussion, we can construct a complete Eagle encoding and decoding algorithm. The entire algorithm consists of three processes: generating parameters, encoding, and decoding.
[Parameter generation]
Firstly we choose two $L$-bit parameters $w_{0}$ and $w_{1}$ which have odd number of different bits, then we choose $L$-bit initial state $s_{0}$.
[Encoding]
For input data $m$, we record $m[i](1\leq i\leq L)$ as the $i$-th bit of $m$, $m[i]\in\{0,1\}$, $L$ is the length of $m$, the encoding process is as follows.
Algorithm 1 EagleEagle-Encoding.
0:Parameter: $(w_{0},w_{1})$. Input: m=[m[1],m[2],,m[L]]m=[m[1],m[2],...,m[L]]. Random number: $s_{0}$.
0:Output: $(c,s_{L})$,
1: Define $c$ as a bit array with length $L$.
2:for $i=1;i\leq L;i++$ do
3:  $c[i]\leftarrow s_{i-1}[L]$
4:  if m[i] = 0 then
5:   $s_{i}\leftarrow w_{0}\oplus(s_{i-1}\oplus s_{i-1}^{+1})$
6:  else
7:   $s_{i}\leftarrow w_{1}\oplus(s_{i-1}\oplus s_{i-1}^{+1})$
8:  end if
9:end for
Where the $i$-th bit of $c$ takes the last bit of state $s_{i-1}$.
[Decoding]
The output $(c,s_{L})$ of the above encoding process is used as the input of the decoding process, the decoding process is as follows.
Algorithm 2 EagleEagle-Decoding.
0:Parameter: $(w_{0},w_{1})$. Input: $c=[c[1],c[2],...,c[L]],s_{L}$.
0:Output: $m$.
1:for $i=L;i\geq 1;i$-- do
2:  Define $x$ as a bit. (x=0orx=1)(x=0\,or\,x=1)
3:  Do trial-and-error testing with $s_{i}\oplus w_{0}$ or $s_{i}\oplus w_{1}$ , find the unique $w_{x}(x=0\,or\,x=1)$ satisfy $s_{x}\oplus s_{x}^{+1}=s_{i}\oplus w_{x}$.
4:  $m[i]\leftarrow x$
5:  For the two possible $s_{x}$ satisfy $s_{x}\oplus s_{x}^{+1}=s_{i}\oplus w_{x}$, we select the one which the last bit is equal to $c[i]$ as $s_{i-1}$.
6:end for
Now we give an example of the above algorithms.
[Example 1] Eagle encoding
We choose the parameters as $w_{0}=10010011$, $w_{1}=11000111$, $s_{0}=01011001$, $m=10010101$.
Since $m[1]=1$, we set $w=w_{1}$, then we have
$s_{1}=w\oplus(s_{0}\oplus s_{0}^{+1})=11000111\oplus 11101011=00101100$
Since $m[2]=0$, we set $w=w_{0}$, then we have
$s_{2}=w\oplus(s_{1}\oplus s_{1}^{+1})=10010011\oplus 01110100=11100111$
Since $m[3]=0$, we set $w=w_{0}$, then we have
$s_{3}=w\oplus(s_{2}\oplus s_{2}^{+1})=10010011\oplus 00101000=10111011$
Since $m[4]=1$, we set $w=w_{1}$, then we have
$s_{4}=w\oplus(s_{3}\oplus s_{3}^{+1})=11000111\oplus 11001100=00001011$
Since $m[5]=0$, we set $w=w_{0}$, then we have
$s_{5}=w\oplus(s_{4}\oplus s_{4}^{+1})=10010011\oplus 00011101=10001110$
Since $m[6]=1$, we set $w=w_{1}$, then we have
$s_{6}=w\oplus(s_{5}\oplus s_{5}^{+1})=11000111\oplus 10010011=01010100$
Since $m[7]=0$, we set $w=w_{0}$, then we have
$s_{7}=w\oplus(s_{6}\oplus s_{6}^{+1})=10010011\oplus 11111100=01101111$
Since $m[8]=1$, we set $w=w_{1}$, then we have
$s_{8}=w\oplus(s_{7}\oplus s_{7}^{+1})=11000111\oplus 10110001=01110110$
Take the last bit of $s_{0},s_{1},s_{2},s_{3},...,s_{7}$ as $c=10111001$.
We use $(c,s_{8})=(10111001,01110110)$ as the results generated by Eagle encoding.
Next, we will provide the example of Eagle decoding.
[Example 2] Eagle decoding
Known $w_{0}=10010011$, $w_{1}=11000111$, $s_{8}=01110110$, $c=10111001$, how to compute $m$?
Firstly by $s_{8}=w\oplus(s_{7}\oplus s_{7}^{+1})=>s_{7}\oplus s_{7}^{+1}=s8\oplus w$, we don’t know whether $w=w_{0}$ (the 8-th bit of $m$ is 0) or $w=w_{1}$ (the 8-th bit of $m$ is 1), we need a trial-and-error test.
Let’s assume $w=w_{0}$, we have
$s_{7}\oplus s_{7}^{+1}=s_{8}\oplus w_{0}=01110110\oplus 10010011=11100101$
There is no such $s_{7}$ satisfies $s_{7}\oplus s_{7}^{+1}=11100101$.
[…]
Let’s assume $w=w_{1}$, we have
$s_{7}\oplus s_{7}^{+1}=s_{8}\oplus w_{1}=01110110\oplus 11000111=10110001$
There exists two solutions $s_{7}=01101111$ and $s_{7}=10010000$, these two solutions are bit reversed.
Since the 8-th bit of $c$ is $c[8]=1$, the last bit of $s_{7}$ is 1, we have $s_{7}=01101111$.
According to the above, because $w=w_{1}$, the 8-bit of $m$ is 1.
Similarily we get $m=10010101$.
It is not difficult to find that the above encoding process and decoding process are correct, that is $(c,s_{L})$ generated by encoding from $m$ can be completely restored through the decoding process. In addition, the encoding process is sequential encoding in the order of $m$’s bits, and the decoding process is sequential decoding in the reverse order of $c$’s bits.
$s_{0}$$s_{1}$$s_{2}$$s_{L}$$s_{0}$$s_{1}$$s_{2}$$s_{L}$$\scriptstyle{m[1]}$$\scriptstyle{m[2]}$$\scriptstyle{m[3]}$$\scriptstyle{m[L]}$$\scriptstyle{c[1]}$$\scriptstyle{c[2]}$$\scriptstyle{c[3]}$$\scriptstyle{c[L]}$
We also noticed the fact that in the above encoding and decoding process, all inputs and outputs do not need to appear $s_{0}$ , This means that the selection of $s_{0}$ will not affect the correctness of the encoding process and decoding process. The arbitrary of $s_{0}$ will bring the unpredictability of the encoded output.
For the convenience of the discussion in the following chapters, here we briefly analyze the effect of unpredictability of $s_{0}$ on the encoded output.
Given the parameters $w_{0}$ and $w_{1}$ that have odd number of different bits, for a certain input $m$ of $L$ bits, since $s_{0}$ is arbitrarily selected, it is obvious that $c$ is uncertain, but is the final state $s_{L}$ necessarily uncertain?
In fact, the answer is no. In some cases, such as $L=2^{u}$ (that is, the parameter length is the power of 2), the final state $s_{L}$ is determined for different choices of $s_{0}$. The final state variable $s_{L}$ which is the output of the encoding process is only related to the input $m$ and has nothing to do with the choice of the initial state $s_{0}$. See Theorem 2 for details.
[Theorem 2] In Eagle encoding algorithm, given the parameters $w_{0}$ and $w_{1}$ that have $L$ bits with odd bits different, for a certain $L$ bit input $m$, if $L=2^{u}$ is satisfied, then for any initial state $s_{0}$, after the Eagle encoding process, the final state $s_{L}$ is only related to the input $m$, and is unrelated with the choice of the initial state $s_{0}$.
Proof.
We represent $m$ as binary stream $x_{1}x_{2}...x_{L}$ , which $x_{i}\in\{0,1\}$, $1\leq i\leq L$.
We execute the Eagle encoding process to $m$ from $x_{1}$ to $x_{L}$ as follows.
$s_{1}=w_{x_{1}}\oplus(s_{0}\oplus s_{0}^{+1})=f_{1}(w_{x_{1}})\oplus(s_{0}\oplus s_{0}^{+1})$
$s_{2}=w_{x_{2}}\oplus(s_{1}\oplus s_{1}^{+1})=f_{2}(w_{x_{1}},w_{x_{2}})\oplus(s_{0}\oplus s_{0}^{+2})$
$s_{3}=w_{x_{3}}\oplus(s_{2}\oplus s_{2}^{+1})=f_{3}(w_{x_{1}},w_{x_{2}},w_{x_{3}})\oplus(s_{0}\oplus s_{0}^{+1}\oplus s_{0}^{+2}\oplus s_{0}^{+3})$
$s_{4}=w_{x_{4}}\oplus(s_{3}\oplus s_{3}^{+1})=f_{4}(w_{x_{1}},w_{x_{2}},w_{x_{3}},w_{x_{4}})\oplus(s_{0}\oplus s_{0}^{+4})$
It is not difficult to find that for any $m=2^{v}$, $s_{m}=f_{m}(w_{x_{1}},...,w_{x_{m}})\oplus(s_{0}\oplus s_{0}^{+m})$ holds, this can be proved by a simple mathematical induction.
In fact, the conclusion is correct for $v=1$.
We assume that the conclusion is correct for v1v-1 , we have
$s_{m/2}=f_{m/2}(w_{x_{1}},...,w_{x_{m/2}})\oplus(s_{0}\oplus s_{0}^{+m/2})$
Since $s_{m/2}$ to $s_{m}$ must do calculations with $m/2$ steps, we have
$s_{m}=f_{m}(w_{x_{1}},.,w_{x_{m}})\oplus(s_{0}\oplus s_{0}^{+m/2})\oplus(s_{0}^{+m/2}\oplus s_{0}^{+m})$
$=f_{m}(w_{x_{1}},.,w_{x_{m}})\oplus(s_{0}\oplus s_{0}^{+m})$
[…]
From theorem 2, for any parameters $w_{0}$ and $w_{1}$ with length $L=2^{u}$, for any initial state $s_{0}$, execute Eagle encoding on $m$ to obtain $s_{L}$ which is irrelevant with $s_{0}$.
Next we introduce two symbols $\mathbb{E}$ and $\mathbb{D}$.
$\mathbb{E}_{w_{0},w_{1}}(m,s_{0})=(c,s_{L})$ : use the key $(w_{0},w_{1})$ to execute Eagle encoding algorithm (Algorithm 1) on any initial state $s_{0}$ and input $m$ to obtain $c$ and $s_{L}$.
$\mathbb{D}_{w_{0},w_{1}}(c,s_{L})=(m,s_{0})$ : use the key $(w_{0},w_{1})$ to execute Eagle decoding algorithm (Algorithm 2) on any $c$ and $s_{L}$ to obtain $s_{0}$ and $m$.
$\mathbb{E}$ and $\mathbb{D}$ both represent a complete Eagle encoding process and Eagle decoding process, and their introduction is mainly for the convenience of deriving encryption algorithms later.
In all the following chapters of this paper, we assume the length $L$ is a power of 2.
Next, we introduce another useful conclusion about the Eagle encoding algorithm.
[Theorem 3] For any equation $\mathbb{E}_{w_{0},w_{1}}(m,*)=(c,*)$, which $(w_{0},w_{1})$ are unknown variables and $m,c$ are known variables, this equation can be seen as an identity, valid for all $(w_{0},w_{1})$.
Proof.
This theorem requires us to construct an algorithm that, given any $w_{0},w_{1}$ to find a valid equation $\mathbb{E}_{w_{0},w_{1}}(m,*)=(c,*)$.
When given $w_{0},w_{1}$, randomly choose $s_{0}^{{}^{\prime}}$, excute $\mathbb{E}_{w_{0},w_{1}}(m,s_{0}^{{}^{\prime}})=(c^{{}^{\prime}},s_{L})$, according to the theorem 2, $s_{L}$ is irrelevant to $s_{0}^{{}^{\prime}}$, then we excute $\mathbb{D}_{w_{0},w_{1}}(c,s_{L})=(m,s_{0})$, thus $\mathbb{E}_{w_{0},w_{1}}(m,s_{0})=(c,s_{L})$. That is to say, when only $m,c$ is known, any $w_{0},w_{1}$ is valid.
If we construct an encryption algorithm where $(w_{0},w_{1})$ is the key, $m$ is the plaintext, $c$ is the ciphertext, $s_{0}$ and $s_{L}$ never appear in the plaintext sequences or the ciphertext sequences, according to the conclusion of Theorem 3, for a cracker, under the condition of only knowing the plaintext ciphertext correspondence $m,c$, any value $(w_{0},w_{1})$ in the key space is valid and is indistinguishable, this feature allows us to construct the security that meets the requirements of computational complexity.
4 Construction of one-way functions
4.1 Introduction to one-way functions
Before constructing the one-way function, we briefly introduce the properties of one-way function and the relationships with the P vs NP problem.
[Definition 1] A function is a one-way function means that the function satisfies the following properties:
a) For a given $x$, there exists a polynomial-time algorithm that output $f(x)$.
b) Given $y$, it is difficult to find an $x$ that satisfies $y=f(x)$, that is, there does not exists a polynomial-time algorithm that finding the $x$.
The NP-complete problem refers to a set of problems that are verifiable in polynomial-time algorithm. For all NP-complete problems, whether there exists algorithms that are solvable in polynomial-time, this is the P vs NP problem. If $P\neq NP$, then for some NP problems, there is no algorithm that is solvable in polynomial-time. If $P=NP$, then for all NP problems, there exists algorithms that are solvable in polynomial-time.
If one-way function exists, it means that there exists such an NP problem, which has no deterministic polynomial time solvable algorithm, that is, $P\neq NP$. This is a direct inference, which can be directly described as the following theorem, See [9] for details.
[Theorem 4] If one-way function exists, then $P\neq NP$.
4.2 Construction of one-way functions
For short key encryption algorithms, the security of the key depends on the computational complexity of cracking the key when given the known plaintext-ciphertext pairs. In theory, if the key can only be cracked through exhaustive search, this algorithm is considered computationally secure. However, currently all short key encryption algorithms, with known plaintext-ciphertext pairs, have no evidence to suggest that attackers can crack the key through only exhaustive search.
In this chapter, we constructed a short key symmetric encryption algorithm, which we named it Eagle. The key consists of four independent variables, each of which contains at least $2^{L-1}$ possible values. During the encryption process, several intermediate random numbers are generated at the encryption party, after the encryption process is completed, these intermediate random numbers are directly deleted at the encryption party, and only the ciphertext is sent to the decryption party. Under the condition of knowing the correspondence between any multiple sets of plaintext-ciphertext, the problem of cracking the key is equivalent to solving a system of equations. In the process of solving this system of equations, at least one free unkown variables cannot be eliminated, so it is necessary to traverse all possible values of at least this variable, which means that the computational complexity for cracking the key is at least $O(2^{L})$.
On the other hand, when the key known, both the encryption and decryption processes can be completed in linear time to the length of the key, meaning that the computational complexity for both encryption and decryption is $O(KL)$, where $K$ is a constant.
Furthermore, since the decryption process under the condition of knowing the ciphertext and key is reversible for the problem of cracking the key under the condition of knowing the correspondence between plaintext and ciphertext, the decryption process can be completed in $O(KL)$, while the cracking process is completed at least in $O(2^{L})$, thus satisfying the condition of one-way function.
It should be noted that although the computational complexity of cracking the key is exponentially under the condition of knowing the correspondence between plaintext and ciphertext, this does not mean that this encryption algorithm is safe and practical. Assume the length of the key is 4L24L-2 bits, although the computational complexity of cracking the key is not less than $O(2^{L})$, may be much lower than the expected complexity of $O(2^{4L-2})$. A secure and practical encryption algorithm needs to ensure that the cracking key can only exhaustively search for every possible key which is the issue to be considered in the next chapter. In this chapter, our goal is to design a one-way function where we only need to design and prove that the computational complexity for cracking keys is exponential to the length of the key.
Next, we will describe it in five sections. Section 4.2.1 introduces the key generation of the Eagle encryption algorithm, Section 4.2.2 introduces the process of the Eagle encryption algorithm, Section 4.2.3 introduces the process of the Eagle decryption algorithm, Section 4.2.4 analyzes the correctness and complexity of the Eagle encryption algorithm, and Section 4.2.5 provides the security analysis of the Eagle algorithm which give a construction of one-way function.
4.2.1 $Eagle$ key generator
The key is $(w_{0},w_{1})$ and $(w_{2},w_{3})$, where $w_{0}$ and $w_{1}$ have odd number of bits different, $w_{2}$ and $w_{3}$ have odd number of bits different.
Assuming $w_{0},w_{1},w_{2},w_{3}$ are all $L$ bits, the length of the entire key can be considered as 4L24L-2 bits, this is because they have a loss of 2 bit.
We randomly generate a $2L$ bit number, taking the first $L$ bit as $w_{0}$. When the last $L$ bit has an odd number of bits that are different from $w_{0}$, take the last $L$ bit data as $w_{1}$. When the last $L$ bit has an even number of bits that are different from $w_{0}$, take the last bit inverted to the last $L$ bit as $w_{1}$.
For example, we generate a number with 16 bits as $1001001101001000$, we set $w_{0}=10010011$ (the first 8 bits) and $w_{1}=01001000$ (the last 8 bits), when $w_{0}$ and $w_{1}$ have 6 (even) bits different, we set the last bit of $w_{1}$ inverted as $w_{1}=01001001$.
For $(w_{0},w_{1})$, because the probability of odd bits different are the same as even bits different, the total number of $(w_{0},w_{1})$ is $2^{2L-1}$.
For $(w_{2},w_{3})$, the generation method is the same as $(w_{0},w_{1})$, the total number of $(w_{2},w_{3})$ is also $2^{2L-1}$.
4.2.2 $Eagle$ encryption process
The plaintext $P$ is grouped by L1L-1 bits, and the last group with less than L1L-1 bits are randomly filled into L1L-1 bits. The total number of groups is assumed to be $K$, denoted as $P=(P_{1},P_{2},...,P_{K})$.
Then, one bit is randomly filled into the last bit of each $P_{i}(1\leq i\leq K)$, making the length of $P_{i}$ equal to $L$.
Generate random numbers of $K$ groups with $L$ bits completely independently, denoted as $(x_{1},x_{2},...,x_{K})$.
Generate another three random numbers denoted as $S_{0}$, $S_{0}^{{}^{\prime}}$ and $y_{1}$.
The encryption process is described in the following algorithm.
Algorithm 3 EagleEagle-Encrypt.
0:Key: $(w_{0},w_{1})$, $(w_{2},w_{3})$. Plaintext sequence: $P=(P_{1},P_{2},...,P_{K})$. Random number sequence: $(x_{1},x_{2},...,x_{K})$. Random number: $S_{0},S_{0}^{{}^{\prime}},y_{1}$.
0:Ciphertext sequence: $(C_{1},C_{2},...,C_{2K})$ and $(C_{1}^{{}^{\prime}},C_{2}^{{}^{\prime}},S_{2}^{{}^{\prime}})$.
[…]
2:for $i=1;i\leq K;i++$ do
3:  $(C_{2i-1},S)\leftarrow\mathbb{E}_{w_{0},w_{1}}(P_{i},S)$
4:  $S\leftarrow(S\oplus(1\&x_{i}))$ {The last bit of $S$ needs to be XORed with the last bit of $x_{i}$ to hide the last bit of $S$ from $C_{2i}$}
5:  $(C_{2i},S)\leftarrow\mathbb{E}_{w_{0},w_{1}}(x_{i},S)$
6:  if $i<K$ then
7:   $S\leftarrow(S\oplus(1\&P_{i+1}))$ {The last bit of $S$ needs to be XORed with the last bit of $P_{i+1}$ to hide the last bit of $S$ from $C_{2i+1}$}
8:  end if
9:end for
[…]
The above algorithm can be represented by the following figure.
$S_{0}$$C_{1},S_{1}$$C_{2},S_{2}$$w_{0},w_{1}$$C_{2},S_{2}$$C_{3},S_{3}$$C_{4},S_{4}$$C_{4},S_{4}$$C_{5},S_{5}$$C_{6},S_{6}$$w_{2},w_{3}$$S_{0}^{{}^{\prime}}$$C_{1}^{{}^{\prime}},S_{1}^{{}^{\prime}}$$C_{2}^{{}^{\prime}},S_{2}^{{}^{\prime}}$$\scriptstyle{P_{1}}$$\scriptstyle{x_{1}}$$\scriptstyle{P_{2}}$$\scriptstyle{x_{2}}$$\scriptstyle{P_{3}}$$\scriptstyle{x_{3}}$$\scriptstyle{y_{0}=S_{6}}$$\scriptstyle{y_{1}}$
4.2.3 $Eagle$ decryption process
For the decrypting party, the key is also $(w_{0},w_{1}),(w_{2},w_{3})$, the ciphertext is $(C_{1}$, $C_{2}$, …, $C_{2K})$ and $(C_{1}^{{}^{\prime}},C_{2}^{{}^{\prime}},S_{2}^{{}^{\prime}})$. Please refer to the following for details.
Algorithm 4 EagleEagle-Decrypt.
0:Key: $(w_{0},w_{1}),(w_{2},w_{3})$. Ciphertext sequence: $(C_{1},C_{2},...,C_{2K})$ and $(C_{1}^{{}^{\prime}},C_{2}^{{}^{\prime}},S_{2}^{{}^{\prime}})$.
0:Plaintext sequence: $P=(P_{1},P_{2},...,P_{K})$.
[…]
6:for $i=K;i\geq 1;i$-- do
7:  $(x_{i},S)\leftarrow\mathbb{D}_{w_{0},w_{1}}(C_{2i},S)$
8:  $S\leftarrow(S\oplus(1\&x_{i}))$
9:  $(P_{i},S)\leftarrow\mathbb{D}_{w_{0},w_{1}}(C_{2i-1},S)$
10:  $S\leftarrow(S\oplus(1\&P_{i}))$
11:end for
4.2.4 Basic analysis of EagleEagle
It is obvious that $Eagle$ encryption algorithm is correct because the decryption process and encryption process are mutually inverse.
From the encryption process above, it can be seen that the $Eagle$ encryption algorithm is divided into two stages: the preparation stage and the encryption stage. In the preparation stage, different parameters are randomly generated. In the encryption stage, the ciphertexts generated by the random parameters and the plaintexts together with the keys. After the encryption process is completed, these independent random parameters are directly deleted by the encryption party and do not appear in the ciphertexts.
It should also be noted that after being encrypted by the $Eagle$ encryption algorithm, the length of the ciphertext is about 2 times the length of the plaintext. This is because in the preparation stage of the $Eagle$ encryption algorithm, independent random numbers with a length of about 1 times the plaintext are generated, and then these random numbers are encrypted using the known short key. The encrypted ciphertext formed by these random numbers is also bound to the final ciphertexts.
The $Eagle$ encryption algorithm performs bit by bit in encryption and decryption process, and the computational complexity of operation for each bit is also a constant level. Therefore, the encryption and decryption complexity of the $Eagle$ encryption algorithm can be regarded as $O(KL)$.
4.2.5 A construction of one-way function
When given the ciphertext ($C_{1}$, $C_{2}$, …, $C_{2K}$) and $(C_{1}^{{}^{\prime}},C_{2}^{{}^{\prime}},S_{2}^{{}^{\prime}})$ of the $Eagle$ encryption algorithm, we define the following function.
$f(w_{0},w_{1},w_{2},w_{3})=(P_{1},P_{2},...,P_{K})$ (4.1)
where $(w_{0},w_{1},w_{2},w_{3})$ is the short key and $(P_{1},P_{2},...,P_{K})$ is the plaintext decrypted by the $Eagle$ decryption algorithm.
The inverse function of $f$ can be defined as
$f^{-1}(P_{1},P_{2},.,P_{K})=\{(w_{0},w_{1},w_{2},w_{3})|$ (4.2)
$f(w_{0},w_{1},w_{2},w_{3})=(P_{1},P_{2},.,P_{K})\}$
Solving $f^{-1}$ is equivalent to finding the key $(w_{0},w_{1},w_{2},w_{3})$ to match the plaintext $(P_{1},P_{2},...,P_{K})$.
In fact, if we can prove that we can’t solve $f^{-1}$ in polynomial time, that is, for any attacker who only knows the plaintext-ciphertext pairs, there is no effective method in polynomial time to find the correct key.
[…]
Intuitively, the Eagle encryption algorithm is composed of two independent parts, each consisting of its own key, plaintext, and ciphertext. Given the correspondence between the plaintext ($P_{1},P_{2},...,P_{K}$) and ciphertext ($C_{1}$, $C_{2}$, …, $C_{2K}$), $(C_{1}^{{}^{\prime}},C_{2}^{{}^{\prime}},S_{2}^{{}^{\prime}})$ of the Eagle encryption algorithm, the problem of solving the key $(w_{0}$, $w_{1}$, $w_{2}$, $w_{3})$ is equivalent to the problem of solving the respective keys in two independent encryption algorithms under the condition that only their plaintext or ciphertext is known. The first encryption algorithm solves its key $(w_{0},w_{1})$ under the condition that its plaintext $(P_{1},P_{2},...,P_{K})$ and its ciphertext $(C_{1},C_{2},...,C_{2K})$ are known. And in another encryption algorithm, the key $(w_{2},w_{3})$ need to be solved under the condition that only the ciphertext ($C_{1}^{{}^{\prime}}$, $C_{2}^{{}^{\prime}}$, $S_{2}^{{}^{\prime}}$) is known.
In fact, for these two independent encryption algorithms, under the condition of only knowing their respective plaintext or ciphertext, every possible key cannot be reasonably excluded, and every possible key must be exhaustively searched. From the perspective of the entire Eagle encryption algorithm, at least one set of keys from $(w_{0}$, $w_{1})$ or $(w_{2}$, $w_{3})$ needs to be exhaustively searched to find a valid key. The computational complexity of exhausting either $(w_{0}$, $w_{1})$ or $(w_{2}$, $w_{3})$ is exponentially.
Next, let’s roughly introduce the idea of the proof. During the encryption process, all generated intermediate states are unknown in the ciphertext, and these intermediate states need to be gradually restored during the decryption process
Without specifying any intermediate state $S$ or $y_{0}$, the problem of solving $f^{-1}$ is equivalently transformed into solving a system of equations about the unknown variables as $(w_{0}$, $w_{1}$, $w_{2}$, $w_{3}$) and $(x_{1},x_{2},...,x_{K})$ and $y_{1}$, Any equation in this system of equations can be regarded as an identity equation about the unknown variables, and it is impossible to find an effective solution.
On the other hand, given $(P_{1}$, $P_{2}$, … , $P_{K})$ and $(C_{1}$, $C_{2}$, …, $C_{2K})$ and $(C_{1}^{{}^{\prime}},C_{2}^{{}^{\prime}},S_{2}^{{}^{\prime}})$, every intermediate state $S$ can take any value, the computational complexity of traversing every intermediate state is $O(2^{L})$.
This indicates that the computational complexity of solving $f^{-1}$ is at least $O(2^{L})$.
Let’s give the following theorem.
[Theorem 5] When given the ciphertext ($C_{1}$, $C_{2}$, …, $C_{2K})$, $(C_{1}^{{}^{\prime}},C_{2}^{{}^{\prime}},S_{2}^{{}^{\prime}}$) and the plaintext $(P_{1},P_{2},...,P_{K})$ of the $Eagle$ encryption algorithm, the computational complexity of the problem of cracking the key $(w_{0}$, $w_{1}$, $w_{2}$, $w_{3})$ is at least $O(2^{L})$.
Proof.
The entire proof process is divided into the following two parts.
(i) Any intermediate state $S$ is unknown
By observing the structure of Algorithm 3 and Algorithm 4, without knowing any $S$, the problem of cracking the key $(w_{0}$, $w_{1}$, $w_{2}$, $w_{3})$ is equivalent to solving the following system of equations
[…]
Where * are placeholder. $(P_{1}$, $P_{2}$, …, $P_{K})$, $(C_{1}$, $C_{2}$, …, $C_{2K})$, $(C_{1}^{{}^{\prime}}$, $C_{2}^{{}^{\prime}}$, $S_{2}^{{}^{\prime}})$ are known variables. $(w_{0},w_{1},w_{2},w_{3})$, $(x_{1},x_{2},...,x_{K})$, $y_{1}$ are unknown variables to be solved.
According to the Theorem 3, any variable in any equation of the system of equations (4.3) and (4.4) can take any value, so each equation can be regarded as an identity equation.
For the system of equations (4.5), the first equation is also an identity. The unknown variable $y_{1}$ in the second equation is a free variable, this equation holds for any $(w_{2},w_{3})$, thus we need to traverse all $(w_{2},w_{3})$ to solve this equation.
We can conclude that the computational complexity of the problem of cracking the key $(w_{0}$, $w_{1}$, $w_{2}$, $w_{3})$ is at least $O(2^{2L-1})$.
(ii) One of the intermediate states $S$ is known.
Without loss of generality we suppose $S_{1}$ is known,according to the Theorem 2, $S_{1}$ depends only on $P_{1}$ and $(w_{0},w_{1})$. For any given combination of $S_{1}$ and $P_{1}$, we can always find some reasonable $(w_{0},w_{1})$, which requires a conclusion that we haven’t yet demonstrated, we will provide a detailed argument in the next chapter, see Theorem 7 for details.
Since we can find some reasonable $(w_{0},w_{1})$ for any given $S_{1}$, thus we need to traverse all $S_{1}$ to solve $(w_{0},w_{1})$ and $(w_{2},w_{3})$.
We can conclude that the computational complexity of the problem of cracking the key $(w_{0}$, $w_{1}$, $w_{2}$, $w_{3})$ is at least $O(2^{L})$.
According to the conclusion of Theorem 5, the computational complexity of solving $f^{-1}$ is exponential.
Due to the fact that solving $f$ can be completed in polynomial time and the computational complexity of solving $f^{-1}$ is exponential, $f$ is a one-way function, we can summarize it to the following theorem.
[Theorem 6] The function $f$ defined in (4.1) is a one-way function.
[…]
5 How to achieve only exhaustive search to crack keys
5.1 In depth analysis of Eagle encoding algorithm
Back to the Eagle encoding algorithm, we noticed that $w_{0}$ differs from $w_{1}$ with an odd number of bits, which is equivalent to the binary representation of $w_{0}\oplus w_{1}$ containing an odd number of 1. There are $2^{2L-1}$ pairs of $(w_{0},w_{1})$ that satisfy an odd number of different bits, This is because among all $(w_{0},w_{1})$, $w_{0}$ and $w_{1}$ contain the same number of odd bits different as even bits different.
Next, let’s take a closer look at the features of the Eagle encoding algorithm. For the parameter $(w_{0},w_{1})$, we have another completely equivalent representation $(\delta,w_{0})$, where $\delta=w_{0}\oplus w_{1}$, which means that $(w_{0},w_{1})$ and $(\delta,w_{0})$ can be equivalently transformed. Since the selection of $s_{0}$ is independent of $s_{L}$, without loss of generality, we set $s_{0}=0$, when given $(\delta,w_{0})$, we only observe the relationship between the final state $s_{L}$ and the input $m=b_{1}b_{2}...b_{L}$, where $b_{i}\in\{0,1\}$. We excute the Eagle encoding algorithm with the first bit to obtain
$s_{1}=w_{b_{1}}\oplus(s_{0}\oplus s_{0}^{+1})=w_{0}\oplus g_{1}(b_{1},\delta)$
Where $g_{1}(b_{1},\delta)=0$ when $b_{1}=0$ and $g_{1}(b_{1},\delta)=\delta$ when $b_{1}=1$. We excute the Eagle encoding algorithm with the second bit to obtain
$s_{2}=w_{b_{2}}\oplus(s_{1}\oplus s_{1}^{+1})$
$=(w_{0}\oplus(w_{0}\oplus w_{0}^{+1}))\oplus g_{2}(b_{1},b_{2},\delta)$
$=w_{0}^{+1}\oplus g_{2}(b_{1},b_{2},\delta)$
Where $g_{2}(b_{1},b_{2},\delta)$ is a function about $\delta$ when $b_{1},b_{2}$ is given. We excute the Eagle encoding algorithm with the third bit to obtain
$s_{3}=w_{b_{3}}\oplus(s_{2}\oplus s_{2}^{+1})$
$=(w_{0}\oplus w_{0}^{+1}\oplus w_{0}^{+2})\oplus g_{3}(b_{1},b_{2},b_{3},\delta)$
So far we see that $s_{i}=h_{i}(w_{0})\oplus g_{i}(b_{1},b_{2},...,b_{i},\delta)$, where $h_{i}(w_{0})$ can be regarded as the state obtained by executing $i$ bits Eagle encoding algorithm when $m=0$, $g_{i}(b_{1},b_{2},...,b_{i},\delta)$ is a function about $\delta$ when $b_{1},b_{2},...,b_{i}$ are given.
When $L=2^{u}$, we set $m=0,s_{0}=0$, after executing Eagle encoding, the final state is $s_{L}=w_{0}^{+(L-1)}$. When $m\neq 0$, after excuting Eagle encoding, the final state is
$s_{L}=w_{0}^{+(L-1)}\oplus g_{L}(b_{1},b_{2},...,b_{L},\delta)$ (5.1)
Where $g_{L}(b_{1},b_{2},...,b_{L},\delta)$ can also be expressed as $g_{L}(m,\delta)$. It is not difficult to see that $g_{L}(m,\delta)$ is a series of XOR operations with $\delta^{+k}$ given $m$, this is also equivalent to $g_{L}(m,\delta)=c_{1}\delta^{+1}\oplus c_{2}\delta^{+2}\oplus...\oplus c_{L}\delta^{+L}$, where $c_{i}\in\{0,1\}$ and $(c_{1},c_{2},...,c_{L})$ only determined by $m$. So (5.1) can ultimately be expressed as
$s_{L}=w_{0}^{+(L-1)}\oplus g_{L}(m,\delta)=w_{0}^{+(L-1)}\oplus(c_{1}\delta^{+1}\oplus c_{2}\delta^{+2}\oplus...\oplus c_{L}\delta^{+L})$ (5.2)
Furthermore, we observe several important facts, as shown in the following theorems.
[Theroem 7] $\forall m,s$, there exists $2^{L-1}$ different groups $(w_{0},w_{1})$ satisfy that $\mathbb{E}_{w_{0},w_{1}}(m,*)=(*,s)$. For any $\delta$, there exists a unique $w_{0}$ that $\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m,*)=(*,s)$.
Proof.
From (5.2), we know $\mathbb{E}_{w_{0},w_{1}}(m,*)=(*,w_{0}^{+(L-1)}\oplus g_{L}(m,\delta))$, through a simple transformation we obtain $w_{0}^{+(L-1)}=g_{L}(m,\delta)\oplus s$, it is obvious that for any $\delta,m,s$, the right side of the equation is a difinite result, so there exists a unique $w_{0}$ valid.
Additionally, due to the odd number of bits differ between $w_{0}$ and $w_{1}$, obviously $\delta=w_{0}\oplus w_{1}$ have an odd number of bits that are 1, that is to say, there are $2^{L-1}$ different $\delta$.
The conclusion of Theorem 7 indicates that, for any $\mathbb{E}_{w_{0},w_{1}}(m,*)=(*,s)$, when only given $(m,s)$, there are exponentially different $(w_{0},w_{1})$ valid.
[Theorem 8] For any $m_{1},s_{1},m_{2},s_{2}$, when the parity between $m_{1}$ and $m_{2}$ is different, there exists a unique $(\delta,w_{0})$ that $\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{1},*)=(*,s_{1})$ and $\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{2},*)=(*,s_{2})$.
Proof.
The state obtained by excuting the Eagle encoding algorithm with L1L-1 bits on $m_{1}$ is $s_{L-1,1}=h_{L-1,1}(w_{0})\oplus g_{L-1,1}(m_{1},\delta)$, the state obtained by excuting the Eagle encoding algorithm with L1L-1 bits on $m_{2}$ is $s_{L-1,2}=h_{L-1,2}(w_{0})\oplus g_{L-1,2}(m_{2},\delta)$.
Due to the difference in parity between $m_{1}$ and $m_{2}$, it can be assumed that the last bit of $m_{1}$ is 0 and the last bit of $m_{2}$ is 1. Further execute the Eagle encoding algorithm with one more bit separately to obtain
$s_{1}=s_{L,1}=w_{0}^{+(L-1)}\oplus(g_{L-1,1}(m_{1},\delta)\oplus g_{L-1,1}(m_{1},\delta)^{+1})$
$s_{2}=s_{L,2}=w_{0}^{+(L-1)}\oplus\delta\oplus(g_{L-1,2}(m_{2},\delta)\oplus g_{L-1,2}(m_{2},\delta)^{+1})$
Perform XOR operation on both sides of the above two equations simultaneously, set $(g_{L-1,1}(m_{1},\delta)\oplus(g_{L-1,2}(m_{2},\delta)$ as $g(m_{1},m_{2},\delta)$ to obtain
$s_{1}\oplus s_{2}=\delta\oplus(g(m_{1},m_{2},\delta)\oplus g(m_{1},m_{2},\delta)^{+1})$
Obviously, the left side of the equation is a definite number, on the right side of the equation is the result of performing XOR operation on an odd number of $\delta^{+k}$, It is not difficult to verify that the solution to such an equation exists and is unique.
The conclusion of Theorem 8 indicates that, for any $\mathbb{E}_{w_{0},w_{1}}(m_{1},*)=(*,s_{1})$ and $\mathbb{E}_{w_{0},w_{1}}(m_{2},*)=(*,s_{2})$, when only given $m_{1},s_{1},m_{2},s_{2}$, there is a 50% probability that a unique $w_{0},w_{1}$ valid.
[Theorem 9] For any two groups $m_{1},m_{2}$ and $m_{1}^{{}^{\prime}},m_{2}^{{}^{\prime}}$, if $m_{1}\oplus m_{2}=m_{1}^{{}^{\prime}}\oplus m_{2}^{{}^{\prime}}$, excute the Eagle encoding algorithm as (5.2) on $m_{1},m_{2},m_{1}^{{}^{\prime}},m_{2}^{{}^{\prime}}$ to obtain
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{1},*)=(*,w_{0}^{+(L-1)}\oplus g_{L}(m_{1},\delta))$ (5.3)
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{2},*)=(*,w_{0}^{+(L-1)}\oplus g_{L}(m_{2},\delta))$
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{1}^{{}^{\prime}},*)=(*,w_{0}^{+(L-1)}\oplus g_{L}(m_{1}^{{}^{\prime}},\delta))$
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{2}^{{}^{\prime}},*)=(*,w_{0}^{+(L-1)}\oplus g_{L}(m_{2}^{{}^{\prime}},\delta))$
then $g_{L}(m_{1},\delta)\oplus g_{L}(m_{2},\delta)=g_{L}(m_{1}^{{}^{\prime}},\delta)\oplus g_{L}(m_{2}^{{}^{\prime}},\delta)$ holds.
Proof.
Returning to the calculation process of (5.2), it is not difficult to find that
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m\oplus\Delta m,*)=(*,w_{0}^{+(L-1)}\oplus g_{L}(m,\delta)\oplus j(\delta,\Delta m))$
Where $j(\delta,\Delta m)$ is a function only related to $\delta,\Delta m$. Without loss of generality we set $m_{1}\oplus m_{2}=m_{1}^{{}^{\prime}}\oplus m_{2}^{{}^{\prime}}=\Delta m$, (5.2) can be reorganized as
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{1}^{{}^{\prime}},*)=(*,w_{0}^{+(L-1)}\oplus g_{L}(m_{1},\delta)\oplus j(\delta,\Delta m))$ (5.4)
$\mathbb{E}_{w_{0},\delta\oplus w_{0}}(m_{2}^{{}^{\prime}},*)=(*,w_{0}^{+(L-1)}\oplus g_{L}(m_{2},\delta)\oplus j(\delta,\Delta m))$
By simple calculate we can obtain
$g_{L}(m_{1}^{{}^{\prime}},\delta)\oplus g_{L}(m_{2}^{{}^{\prime}},\delta)$
$=g_{L}(m_{1},\delta)\oplus g_{L}(m_{2},\delta)\oplus(w_{0}^{+(L-1)}\oplus w_{0}^{+(L-1)})\oplus(j(\delta,\Delta m)\oplus j(\delta,\Delta m))$
$=g_{L}(m_{1},\delta)\oplus g_{L}(m_{2},\delta)$
Theorem 9 indicates that, the function $g_{L}(m,\delta)$ satisfies the distributive law for both variables $m$ and $\delta$.
[Theorem 10] For the function $g_{L}(m,\delta)$ in (5.2), fix any $\delta$, where $\delta$ has an odd number of bits that are 1. Fix any $u$, there exists a unique $m$ that $g_{L}(m,\delta)=u$.
Proof.
Fix any $\delta$, set $w_{0}\in\{0,1\}^{L},w_{1}=w_{0}\oplus\delta$, excute Eagle encoding algorithm for $m$, we obtain $\mathbb{E}_{w_{0},w_{1}}(m,*)=(*,s)$, then $u=s\oplus w_{0}^{+L}$, $g_{L}(m,\delta)=u$, thus $m\to u$ is monomorphism.
On the other hand, fix any $\delta$, set $w_{0}\in\{0,1\}^{L},w_{1}=w_{0}\oplus\delta$, excute Eagle decoding algorithm for $s=u\oplus w_{0}^{+L}$, we obtain $\mathbb{D}_{w_{0},w_{1}}(*,s)=(m,*)$, thus $u\to m$ is monomorphism.
Furthermore, since the ranges of $m$ and $u$ are the same, the mapping from $m$ to $s$ is a one-to-one corresponding when $\delta$ is fixed.
The proof process of Theorem 10 actually describes the process of $g_{L}(m,\delta)=u$ and its inverse function $g_{L}^{-1}(u,\delta)$.
5.2 The multiplication operation of binary
According to the conclusion of Theorem 9, the symbol $g_{L}(m,\delta)$ satisfies the distributive law for both $m$ and $\delta$, that is
$g_{L}(m,\delta_{1}\oplus\delta_{2})=g_{L}(m,\delta_{1})\oplus g_{L}(m,\delta_{2})$
$g_{L}(m_{1}\oplus m_{2},\delta)=g_{L}(m_{1},\delta)\oplus g_{L}(m_{2},\delta)$
If we consider the XOR operation $\oplus$ as an addition operation to binary, $g_{L}(m,\delta)$ can be seen as the multiplication operation between $m$ and $\delta$, as follows
$g_{L}(m,\delta)=m*\delta$
It is not difficult to verify that $m*\delta$ does not satisfy the commutative law or the associative law, so $m*\delta$ can only be seen as the right multiplication of $\delta$ by $m$. It is obvious that this representation is well-defined.
The calculation of $m*\delta$ is shown in Algorithm 5 as follows.
Algorithm 5 calculation of $m*\delta$.
0:Input: $m,\delta$.
0:Output: $m*\delta$.
[…]
5:return $s_{L}$.
This paper does not intend to delve into the algebraic properties of the right multiplication operation mentioned above, but only simplifies the representation of $g_{L}(m,\delta)$ using $m*\delta$.
Observing the equation $R*\delta*\delta=S$, whereRR and $S$ are known variables, $\delta$ is the unknown variable to be solved. Since the left side of the equation performs two right multiplications on $\delta$, this equation can be regarded as a “quadratic equation in one variable” about $\delta$. Obviously, this is a nonlinear equation, and we have sufficient reason to make the following assertion.
[Assertion 1] The equation $R*\delta*\delta=S$ about $\delta$ can only be solved by exhaustively search for $\delta$.
Without causing ambiguity, we denote the operation of performing two right multiplications by $\delta$ on $R$ as $R*\delta^{2}$.
Similarly, we can also introduce the division operator //, using the notation $S/\delta=R$ to represent $R*\delta=S$, and using $S/\delta^{2}$ to represent performing two division operations on $S$ with respect to $\delta$. The calculation for division is as follows.
Algorithm 6 calculation of $m/\delta$.
0:Input: $s,\delta$.
0:Output: $s/\delta$.
[…]
5:return $m$.
5.3 Design of encryption algorithm for only exhaustive method to crack the key
Although we found an encryption algorithm in Chapter 4 that, given the known correspondence between plaintext and ciphertext, the computational complexity for cracking the key is at least exponentially to the length of key as $O(2^{L})$, this does not mean that this algorithm is practical. Because the length of the key $(w_{0},w_{1},w_{2},w_{3})$ is 4L24L-2 bits, theoretically, the computational complexity for cracking the key must reach $O(2^{4L-2})$ to recognize its secure.
In this chapter, we utilized the Eagle encoding algorithm and its related features to design another short key symmetric block encryption algorithm, which we named the $Eagle^{*}$ encryption algorithm. The main feature of this algorithm is that it introduce a nonlinear operation, together with the random numbers as designed in Eagle encryption algorithm.
After being encrypted by this encryption algorithm, the problem of cracking the key given that all plaintext ciphertext correspondences are known, is equivalent to the problem of two independent encryption systems cracking their respective keys with only their ciphertext known.
In addition, for exhaustively searching for either of the two keys, solving the other key requires solving a nonlinear equation, which can only be solved by exhaustive search.
The entire $Eagle^{*}$ encryption algorithm is divided into three parts: key generation, encryption algorithm, and decryption algorithm. Next, we will provide a detailed description.
5.3.1 $Eagle^{*}$ key generator
The key is divided into three independent variables, denoted as $\delta_{1}$, $\delta_{2}$ and $\delta_{3}$, where $\delta_{1}$, $\delta_{2}$ and $\delta_{3}$ both have odd number of bits with 1.
Assuming $\delta_{1},\delta_{2},\delta_{3}$ are all $L$ bits, the length of the entire key can be considered as 3L33L-3 bits, this is because each variable has a loss of 1 bit.
5.3.2 $Eagle^{*}$ encryption algorithm
The plaintext $P$ is grouped by $L$ bits, and the last group with less than $L$ bits are randomly filled into $L$ bits. The total number of groups is assumed to be $K$, denoted as $P=(P_{1},P_{2},...,P_{K})$.
We randomly generate $2(K+3)$ groups number of $L$ bits, denoted as $x=(x_{1}$, $x_{2}$, …, $x_{K+3})$, $y=(y_{1}$, $y_{2}$, …, $y_{K+3})$, where $x_{i}$ and $y_{i}$ have odd number of bits different.
Then we randomly generate another four number as $P_{K+1},P_{K+2},P_{K+3}$ and $S_{0}$.
The encryption process is detailed in Algorithm 7 below.
Algorithm 7 $Eagle^{*}$-Encrypt.
0:Keys: $(\delta_{1},\delta_{2},\delta_{3})$. Plaintext sequences: $P=(P_{1},P_{2},...,P_{K})$. Random number sequences: $x=(x_{1},x_{2},...,x_{K+3})$,                $y=(y_{1},y_{2},...,y_{K+3})$.Random number: $P_{K+1},P_{K+2},P_{K+3}$ and $S_{0}$.
0:Ciphertext sequences: $C=(C_{1},C_{2},...,C_{K+3},S_{K+3}$                 $Sx_{1},Sx_{2},...,Sx_{K+3},$                 $Sy_{1},Sy_{2},...,Sy_{K+3})$
1:for $i=1;i\leq K+3;i++$ do
2:  $Sx_{i}\leftarrow x_{i}*\delta_{1}*\delta_{1}$
3:  $Sy_{i}\leftarrow y_{i}*\delta_{2}*\delta_{2}$
4:end for
5:for $j=1;j\leq K+3;j++$ do
6:  $(C_{j},S_{j})\leftarrow\mathbb{E}_{x_{j},y_{j}}(P_{j},S_{j-1})$
7:  if $j<K+3$ then
8:   $S_{j}\leftarrow S_{j}*\delta_{3}*\delta_{3}$
9:  end if
10:end for
From the above algorithm, it can be seen that the $Eagle^{*}$ encryption algorithm is divided into two independent parts, as shown in the following figure.
$S_{0}$$x_{1}*\delta_{1}^{2}=Sx_{1};y_{1}*\delta_{2}^{2}=Sy_{1}$$S_{1}*\delta_{3}^{2}$$C_{1}$$x_{2}*\delta_{1}^{2}=Sx_{2};y_{2}*\delta_{2}^{2}=Sy_{2}$$S_{2}*\delta_{3}^{2}$$C_{2}$$S_{K+2}*\delta_{3}^{2}$$C_{K+2}$$x_{K+3}*\delta_{1}^{2}=Sx_{K+3};y_{K+3}*\delta_{2}^{2}=Sy_{K+3}$$S_{K+3}$$C_{K+3}$$\scriptstyle{P_{1},x_{1},y_{1}}$$\scriptstyle{P_{2},x_{2},y_{2}}\scriptstyle{......}$$\scriptstyle{P_{K+3},x_{K+3},y_{K+3}}$
As shown from the above figure, the ciphertext length is about three times the plaintext length.
5.3.3 $Eagle^{*}$ decryption algorithm
For $Eagle^{*}$ decryptors, the ciphertext sequences are $C=($$C_{1}$, $C_{2}$, …, $C_{K+3}$, $Sx_{1}$, $Sx_{2}$, …, $Sx_{K+3}$, $Sy_{1}$, $Sy_{2}$, …, $Sy_{K+3})$ and $S_{K+3}$. Here, the ciphertext sequences can be divided into three independent parts $(C_{1},C_{2}$, …, $C_{K+3},S_{K+3})$ and $(Sx_{1},Sx_{2}$, …, $Sx_{K+3})$ and $(Sy_{1},Sy_{2}$, …, $Sy_{K+3})$, First we use $(\delta_{1})$ and $(Sx_{1},Sx_{2},...,Sx_{K+3})$ to calculate the random number sequences $x=(x_{1}$, $x_{2}$, …, $x_{K+3})$, then we use $(\delta_{2})$ and $(Sy_{1}$, $Sy_{2}$, …, $Sy_{K+3})$ to calculate the random number sequences $y=(y_{1}$, $y_{2}$, …, $y_{K+3})$, then we use the key $(\delta_{3})$ and the ciphertexts $(C_{1}$, $C_{2}$, …, $C_{K+3}$, $S_{K+3})$ and the random number sequences $x=(x_{1}$, $x_{2}$, …, $x_{K+3})$ and $y=(y_{1}$, $y_{2}$, …, $y_{K+3})$ to decrypt $P=(P_{1}$, $P_{2}$, …, $P_{K})$. The specific process is detailed in the following algorithm.
Algorithm 8 $Eagle^{*}$-Decrypt.
0:Key: $(\delta_{1},\delta_{2},\delta_{3})$. Ciphertext sequences: $C=(C_{1},C_{2},...,C_{K+3},S_{K+3},$                 $Sx_{1},Sx_{2},...,Sx_{K+3},$                 $Sy_{1},Sy_{2},...,Sy_{K+3})$.
0:Plaintext sequences: $P=(P_{1},P_{2},...,P_{K})$.
1:for $i=1;i\leq K+3;i++$ do
2:  $x_{i}\leftarrow Sx_{i}/\delta_{1}/\delta_{1}$
3:  $y_{i}\leftarrow Sy_{i}/\delta_{2}/\delta_{2}$
4:end for
5:for $j=K+3;j\geq 1;j$-- do
6:  $(P_{j},S_{j-1})\leftarrow\mathbb{D}_{x_{j},y_{j}}(C_{j},S_{j})$
7:  $S_{j-1}\leftarrow S_{j-1}/\delta_{3}/\delta_{3}$
8:end for
From the processes of Algorithm 7 and Algorithm 8, it can be seen that the $Eagle^{*}$ encryption and $Eagle^{*}$ decryption are mutually inverse, which means that as an encryption algorithm, $Eagle^{*}$ is correct.
In addition, it is easy to see that the calculations are all completed in linear time, so the computational complexity of Algorithm 7 and Algorithm 8 are all $O(KL)$.
5.3.4 Basic Security Analysis
Before delving into the security of the $Eagle^{*}$ encryption algorithm, we first introduce another temporary encryption algorithms, named A-Encrypt algorithm.
The encryption process of the A-Encrypt algorithm is as follows
Algorithm 9 A-Encrypt.
0:Key: $(\delta)$. Plaintext sequences: $R=(R_{1},R_{2},...,R_{K})$.
0:Ciphertext sequences: $C=(S_{1},S_{2},...,S_{K})$.
1:for $i=1;i\leq K;i++$ do
2:  $S_{i}\leftarrow R_{i}*\delta*\delta$
3:end for
The decryption process of the A-Encrypt algorithm is as follows
Algorithm 10 A-Decrypt.
0:Key: $(\delta)$. Ciphertext sequences: $C=(S_{1},S_{2},...,S_{K})$.
0:Plaintext sequences: $R=(R_{1},R_{2},...,R_{K})$.
1:for $i=1;i\leq K;i++$ do
2:  $R_{i}\leftarrow S_{i}/\delta/\delta$
3:end for
Observing the plaintext and ciphertext structure of the $Eagle^{*}$ encryption algorithm, we can easily find the following conclusions.
[Theorem 11] In the $Eagle^{*}$ encryption algorithm, when know the plaintext $P=(P_{1},P_{2},...,P_{K})$ and the corresponding ciphertext $C=(C_{1},C_{2}$, …, $C_{K+3}$, $S_{K+3}$, $Sx_{1}$, $Sx_{2}$, …, $Sx_{K+3}$, $Sy_{1}$, $Sy_{2}$, …, $Sy_{K+3}$), the problem of cracking its key $(\delta_{1},\delta_{2},\delta_{3})$ is equivalent to cracking two respective keys $\delta_{1}$ and $\delta_{2}$ under the condition of knowing only the ciphertext $(Sx_{1},Sx_{2},...,Sx_{K+3})$ of A-Encrypt algorithm and only the ciphertext $(Sy_{1},Sy_{2},...,Sy_{K+3})$ of another A-Encrypt algorithm.
Proof.
According to the definition of $Eagle^{*}$, when know the plaintext $P=(P_{1},P_{2},...,P_{K})$ and the corresponding ciphertext $C=(C_{1},C_{2}$, …, $C_{K+3}$, $S_{K+3}$, $Sx_{1}$, $Sx_{2}$, …, $Sx_{K+3}$, $Sy_{1}$, $Sy_{2}$, …, $Sy_{K+3}$), the problem of cracking its key $(\delta_{1},\delta_{2},\delta_{3})$ is equivalent to solve the following system of equations.
[…]
Where $x_{1},x_{2},...,x_{K+3},y_{1},y_{2},...,y_{K+3},P_{K+1},P_{K+2},P_{K+3},\delta_{1},\delta_{2}$ are unknown variables to be solved. $P_{1}$, $P_{2}$, …, $P_{K+3}$, $Sx_{1}$, $Sx_{2}$, …, $Sx_{K+3}$, $Sy_{1}$, $Sy_{2}$, …, $Sy_{K+3}$, $S_{K+3}$ are known variables.
According to the conclusion of Theorem 3, each equation in the system of equations (5.5) can be regarded as an identity equation.
The unknown variables $x_{1}$, $x_{2}$, …, $x_{K+3}$ in the system of equations (5.6) can be equivalently regarded as the plaintexts of the A-encryption algorithm, with the known variables $Sx_{1}$, $Sx_{2}$, …, $Sx_{K+3}$ being the ciphertexts of the A-encryption algorithm, and $\delta_{1}$ being the key. Solving the system of equations (5.6) is equivalent to crack the key $\delta_{1}$ under the condition of knowing only the ciphertext $(Sx_{1},Sx_{2},...,Sx_{K+3})$ of A-Encrypt algorithm.
Similarly, solving the system of equations (5.7) is equivalent to crack the key $\delta_{2}$ under the condition of knowing only the ciphertext $(Sy_{1},Sy_{2},...,Sy_{K+3})$ of another A-Encrypt algorithm.
Normally, for a short key block encryption algorithm, when only the ciphertext is known, cracking possible keys can only be done using exhaustive search, but this is not necessarily the case. In fact, if an encryption algorithm only knows the ciphertext and meets the following three conditions, it can only use exhaustive search to crack possible keys.
(i) Given any key, any ciphertext in one group is equally possible.
(ii) In the case of knowing only ciphertext, any key for one group is equally possible.
(iii) In the case of knowing only ciphertext, any plaintext in one group is equally possible.
According to the conclusion of Theorem 7, A-Encrypt obviously satisfies the above three conditions. This means that A-Encrypt can only use exhaustive search to crack the key when only knowing their ciphertexts, and there are no effective algorithms.
Next, consider several possible exhaustive attack methods.
[Case 1] Only enumerate either $\delta_{1}$ or $\delta_{2}$.
It suffices to consider only the case where $\delta_{1}$ is exhausted, $x_{1}$, $x_{2}$, …, $x_{K+3}$ can be directly calculated from the system of equations (5.6), but this does not change the conclusion that each equation in the system of equations (5.5) is an identity equation.
[Case 2] Only enumerate $\delta_{3}$.
When only exhaust $\delta_{3}$, since both $x_{i}$ and $y_{i}$ are unknown, it is impossible to determine $S_{i}$ solely based on $S_{K+3}$, which brings no benefit to solving the system of equations (5.5), (5.6), and (5.7).
[Case 3] Only enumerate $(\delta_{1},\delta_{2})$.
When enumerating over $(\delta_{1},\delta_{2})$, we can directly calcute $x_{1}$, $x_{2}$, …, $x_{K+3}$ from the system of equations (5.6), and directly calculate $y_{1}$, $y_{2}$, …, $y_{K+3}$ from the system of equations (5.7). Substituting these calculated $x_{i}$ and $y_{i}$ into the system of equations (5.5), we can sequentially calculate the hidden variables $S_{1}$, $S_{2}$, …, $S_{K+2}$. Then we can solve the equation $S_{i}*\delta_{3}*\delta_{3}=S_{i+1}$ to obtain $\delta_{3}$, this equation is a nonlinear equation about $\delta_{3}$. As long as Assertion 1 holds, the only way to solve this equation is exhaustively enumerating $\delta_{3}$.
[Case 4] Only enumerate $(\delta_{1},\delta_{3})$ or $(\delta_{2},\delta_{3})$.
It suffices to consider only the case where $(\delta_{1},\delta_{3})$ is exhausted, $x_{1}$, $x_{2}$, …, $x_{K+3}$ can be directly calculated from the system of equations (5.6). Since $y_{i}$ is unknown, it is impossible to determine $S_{i}$ solely based on $x_{i}$ and $S_{K+3}$, which brings no benefit to solving the system of equations (5.5), (5.6), and (5.7).
From the above analysis, it can be seen that for any known plaintext-ciphertext correspondence, the correct key can only be found by exhaustively enumerating all $(\delta_{1},\delta_{2},\delta_{3})$.
Next, we need to consider another possible attack method. The attacker finds the relationship between the ciphertext to be decrypted and the ciphertexts- plaintexts in the sample from a large number of plaintext-ciphertext pairs, and directly decrypts or partially decrypts without cracking the key. The premise for this attack method to be effective is that there is some statistical relationship or some specific linear relationship between the plaintext and the ciphertext.
Due to the complete randomness of $x_{i}$ and $y_{i}$, there is no statistical relationship or any form of mathematical relationship between the plaintext $P_{i}$ and $Sx_{i}$. Similarly, there is no statistical relationship between $P_{i}$ and $Sy_{i}$. Furthermore, according to Theorem 3, $P_{i}$ and $C_{i}$ are also completely unrelated. In fact, if we disregard $S_{i}$ and $\delta_{3}$ (in reality, only $S_{K+3}$ and $\delta_{3}$ are known, all $S_{i}$ are invisible without solving for $x_{i}$ and $y_{i}$), the encryption process of $Eagle^{*}$ is equivalent to one-time pad. An attacker cannot find any patterns between plaintexts and ciphertexts from a large sample of plaintext-ciphertext pairs.
Lastly, it should be noted that during encryption, why do we need to randomly generate three additional sets of plaintexts $(P_{K+1},P_{K+2},P_{K+3})$ after $P_{K}$? This is because there exists an implicit relationship between $S_{K+3}$ and $P_{K+3}$ in the ciphertext, which is $S_{K+3}=P_{K+3}*(x_{K+3}\oplus y_{K+3})\oplus x_{K+3}^{+(L-1)}$. In order to make this implicit relationship invisible, we generated several additional sets of plaintext sequences when designing $Eagle^{*}$. The reason why we chose three sets is that there are three variables in the key, so we opted to generate three additional sets of plaintexts.
5.3.5 Linear attack analysis to $Eagle^{*}$ encryption algorithm
Linear attack is a very effective attack method proposed by M. Matsui[2] at the European Cryptology Conference in 1993. Later, scholars quickly discovered that the linear attacks are applicable to almost all block encryption algorithms, and linear attacks have became the main attacks for block encryption algorithms. Various new attacks based on linear attacks are constantly being proposed.
The core idea of linear attack is to take the nonlinear transformation in the cryptographic algorithm, such as the linear approximation of the S-box, and then extend the linear approximation to the linear approximation of the round function, and then connect these linear approximations to obtain a linear approximation of the entire cryptographic algorithm, finally a large number of known plaintext-ciphertext pairs encrypted with the same key are used to exhaustively obtain the plaintext and even the key.
We have noticed that the reason why linear attacks have become an effective attack for block encryption algorithms is that when the key is known, there is a certain implicit linear relationship between the ciphertext and the plaintext. By analyzing the known plaintext-ciphertext pairs, some effective linear relations can be obtained, and some bits of the key can be guessed.
In the $Eagle^{*}$ encryption algorithm, the entire encryption system is divided into three independent parts, each with its own independent key and parameters. The ciphertext of the encryption system that includes the original plaintext is generated by mixing the original plaintext with random parameters, without knowing the random parameters, ciphertext can be regarded as completely independent of the original plaintext, which means that with random parameter changes, ciphertext can equally probability traverse any text in the ciphertext space. And the two other parts of the ciphertext is completely generated by random parameters, completely unrelated to the original plaintext. This can be expressed as
$Pr(C_{i}=c_{i}|P_{i}=p_{i})=\frac{1}{2^{L}}$
$Pr(Sx_{i}=sx_{i}|P_{i}=p_{i})=\frac{1}{2^{L}}$
$Pr(Sy_{i}=sy_{i}|P_{i}=p_{i})=\frac{1}{2^{L}}$
$Pr(S_{K+3}=s|P_{i}=p_{i})=\frac{1}{2^{L}}$
5.3.6 Differential attack analysis to $Eagle^{*}$ encryption algorithm
Differential attack was proposed by Biham and Shamir[6] in 1993, it is a chosen- plaintext attack. Its core idea is to obtain key information by analyzing specific plaintext and ciphertext differences.
The essence of a differential attack is to track the “difference” of the plaintext pair, where the “difference” is defined by the attacker according to the target, which can be an exclusive XOR operation or other target values.
For the $Eagle^{*}$ encryption algorithm, suppose the differential attacker chooses two specific plaintexts $P_{1}$ and $P_{2}$, their difference is $\Delta$, that is $P_{2}=P_{1}+\Delta$, Since the ciphertexts in the second part and the third part are completely unrelated to the plaintext, only the ciphertext in the first part will be analyzed here, the corresponding ciphertexts are $C_{1}$ and $C_{2}$ , and the difference between the ciphertexts is $\varepsilon$ as $C_{2}=C_{1}+\varepsilon$. Since $C_{1}$ and $C_{2}$ are completely random, it is uncertain whether the difference $\varepsilon$ of the ciphertext is caused by randomness of $(x_{1},y_{1},x_{2},y_{2})$ or the uncertainty of the plaintext $(P_{1},P_{2})$. Furthermore, for any key $(\delta_{1},\delta_{2},\delta_{3})$, $x_{1}*\delta_{1}*\delta_{1}$ and $y_{1}*\delta_{2}*\delta_{2}$ subject to the uniform distribution, which can be denoted as
$Pr(C_{1}=c_{1},.,C_{K+3}=c_{K+3}|$
$(P_{1}=p_{1},.,P_{K}=p_{K},W=(\delta_{1},\delta_{2},\delta_{3}))$
$=1/2^{(K+3)L}$
$Pr(Sx_{1}=sx_{1},.,Sx_{K+3}=sx_{K+3}|$
$(P_{1}=p_{1},.,P_{K}=p_{K},W=(\delta_{1},\delta_{2},\delta_{3}))$
$=1/2^{(K+3)L}$
$Pr(Sy_{1}=sy_{1},.,Sy_{K+3}=sy_{K+3}|$
$(P_{1}=p_{1},.,P_{K}=p_{K},W=(\delta_{1},\delta_{2},\delta_{3}))$
$=1/2^{(K+3)L}$
That is to say, for any specific plaintext $P_{1}$ and $P_{2}$ selected by the attacker, after being encrypted with the same key, the corresponding block ciphertexts $(C_{1},C_{2},...,C_{K+3})$, $(Sx_{1},Sx_{2},...,Sx_{K+3})$ and $(Sy_{1},Sy_{2},...,Sy_{K+3})$ are completely random, and every possible value in the ciphertext space appears with equal probability. The attacker has no way to capture the propagation characteristics of the “difference” in the plaintext.